The product of some number plus 2 and 5 is 11 less than the number times 8. Find the number.
Answer by lordacro(12)

To make an equation for it, you have to carefully look at the statement and make some algebra statements of your own.
"Some number plus two" is (x + 2)
The product of that and 5 is 5*(x+2)
"Is" should hint that there is an equals sign.
So far we have 5*(x + 2) =
"The number times 8" means 8x
So eleven less than that is 8x-11.
Put it all together: 5(x+2) = 8x-11
Distribute the five. You have five times x *and* five times two.
5x + 10 = 8x - 11
Get all your x's to one side by subtracting 5x from both sides.
10 = 3x - 11
Now get your numbers together by adding eleven to both sides
21 = 3x
Divide by three.
7 = x